OpenStudy (anonymous):

A5=A5+(5-1) x 8 x|01|02|03|04|05 y|05|13|21|29|37 The difference between the numbers is 8 order of operations says to subtract the 5-1 in the brackets, than add them with the a5. So 5+4 = 9 and i times that by 8? but how does this add up to the x in the 5 term. did I miss a step?

OpenStudy (anonymous):

order of operations says to multiply before you add. first calculate the (5-1) in brackets, next multiply by 8, and finally add 5. (5-1)=4 4*8=32 5+32=37

OpenStudy (anonymous):

order of operations: 1) parentheses 2) exponents 3) multiplication 4) addition

OpenStudy (anonymous):

Aww thank you! XD I havent done this in sooo long!

OpenStudy (anonymous):

:)